Determine wheather the graphs of the equations are perpendicular.
6x-7y=8
6y-7x=5
Are the graphs of the given equations perpendicular?
No or Yes

Found 2 solutions by swincher4391, stanbon:
Answer by swincher4391(1107)

6x - 7y = 8
-7y = 8-6x
y = 6/7x - 8/7
6y - 7x = 5
6y = 5+7x
y = 5/6 + 7/6x
So we have slopes 6/7 and 7/6.
To be perpendicular, the slopes must be negative reciprocals. That is, when you multiply the slopes together you get -1.
(6/7)*(7/6) = 1.
1 =/= -1
So, they are not perpendicular.

Answer by stanbon(75887)

Determine wheather the graphs of the equations are perpendicular.
6x-7y=8
6y-7x=5
Are the graphs of the given equations perpendicular?
No or Yes
--------------------
Check the slopes:
1st Eq: y = (6/7)x-(8/7)
slope = 6/7
---------------------
2nd Eq: y = (7/6)x+(5/6)
------
Not perpendicular because the product of the
slopes is not -1.
Cheers,
Stan H.